A rectangular box is(3x+7) feet wide, (6x + 3) feet long, and (7x-1)feet high.
Find the volume of this box./
Please help me with this one,I tried to do it.
V=LWH
V=(6x+3)(3x+7)(7x-1)
126x-21
Thank you

Answer by scott8148(5880)

(6x+3)(3x+7) = 6x(3x+7)+3(3x+7) = 18x^2+42x+9x+21 = 18x^2+51x+21

(7x-1)(18x^2+51x+21) = 126x^3+357x^2+147x-18x^2-51x-21 = 126x^3+339x^2+96x-21